# Radiated energy fraction of SPI-induced disruptions at ASDEX Upgrade

### v1

@@ Abstract
Future large tokamaks will operate at high plasma currents and high stored plasma energies. To ensure machine protection in case of a sudden loss of plasma confinement (major disruption), a large fraction of the magnetic and thermal energy must be radiated to reduce thermal loads. The disruption mitigation system for ITER is based on massive material injection in the form of shattered pellet injection (SPI). To support ITER, a versatile SPI system was installed at the tokamak ASDEX Upgrade (AUG). The AUG SPI features three independent pellet generation cells and guide tubes, and each was equipped with different shatter heads for the 2022 experimental campaign. We dedicated over 200 plasma discharges to the study of SPI plasma termination, and in this manuscript report on the results of bolometry (total radiation) analysis. We found, that the amount of neon inside the pellets is the dominant factor determining the radiated energy fraction ($\text{f}_{\text{rad}}$). Large and fast fragments, produced by the $12.5^{\circ}$ rectangular shatter head, lead to somewhat higher values of $\text{f}_{\text{rad}}$ compared to the $25^{\circ}$ circular or rectangular heads. This effect is strongest for neon content of $\lesssim 4\times 10^{20}$ neon atoms injected, where a higher normal velocity component (larger fragments) seems slightly beneficial. While full-sized, 8 mm diameter, pure deuterium (D2) pellets lead to a disruption, the 4 mm or shortened 8 mm pellets of pure D2 did not lead to a disruption. The disruption threshold for pure D2 is found to be around $1\times 10^{22}$ deuterium molecules inside the pellet. While the radiated energy fraction of non-disruptive SPI is below 20%, this is increased to 40% during the TQ and VDE phase of the disruptive injections. For deuterium–neon–mix pellets, $\text{f}_{\text{rad}}$-values of $\leq 90$% are observed, and the curve saturates around 80% already for 10% neon mixed into the 8 mm pellets ($2\times 10^{21}$ neon atoms).
@@ 1 Introduction
Plasma terminating disruptions pose a major challenge for tokamaks with high plasma current and stored energy [1, 2, 3]. An efficient disruption mitigation system (DMS) should address three main challenges: minimise the thermal loads on the plasma facing components (PFCs), minimise eddy and halo currents which cause large mechanical loads, while suppressing the formation of a large runaway electron (RE) beam [1]. In present day tokamaks disruption mitigation is typically realised using massive material injection (MMI), either via massive gas injection (MGI) or shattered pellet injection (SPI). High-Z material inside the injection mix helps spread the energy over a larger surface area via radiation, reducing the localised thermal loads [4, 5]. The mechanical loads can be reduced by tailoring the duration of the current quench (CQ) through the injected material composition [4, 6] or early injections in cases of vertical displacement events (VDEs) [7, 8]. The formation of a RE beam is hoped to be suppressed by increasing the free electron density in the plasma core [9], hence the material assimilation (of low Z material – namely protium/deuterium) is a key factor [10, 11, 12].
Shattered pellet injections were first tested at DIII-D in 2009 [13, 14], and over time more tokamaks, such as JET [14], KSTAR [15, 16], HL-2A [17], J-TEXT [18, 19], and EAST [20, 21] followed.
In late 2021, a highly flexible SPI system was installed at ASDEX Upgrade (AUG) [22, 23, 24] to provide further input for the design and optimisation of the ITER DMS [25, 10] based on the SPI principle [26]. In this paper we discuss the evolution of total radiated energy ($\text{W}_{\text{rad}}$) and the radiated energy fraction ($\text{f}_{\text{rad}}$) as a function of pellet and injection parameters in different plasma scenarios, based on the bolometry analysis of the 2022 AUG SPI experimental campaign.
@@ 2.1 Shattered pellet injector on AUG
The AUG SPI is a triple-barrel system, with three independent guide tubes and three different shatter heads [22, 24]. Three pellets, made of mixtures of deuterium and neon (as well as pure deuterium or pure neon) can be generated and launched simultaneously, potentially allowing the study of multi-injection scenarios. A detailed description of the SPI system is given in the paper by Dibon et al [22] and the laboratory commissioning [^2] Animation video of the laboratory setup available at https://datashare.mpcdf.mpg.de/s/DlMzGcWnZwoHMjq or [24]. by Heinrich et al [24].
@@
Figure 2: Comparison of the spatial distribution of the fragment plumes generated by different shatter heads. In (a) the a larger spread of the fragments is observed for the circular, $25^{\circ}$ shatter head compared to the more collimated fragment plume of the rectangular, $12.5^{\circ}$ head for similar pre-shattering pellet velocity in (b). In (c) and (d) the geometrical effects of the circular shatter head are shown. The effective shatter angle depends on the impact position ($\theta$) inside a circular shatter head [27], while for the rectangular head the shatter angle is equal to the miter bend angle $\alpha$. Additionally, the shards may receive perpendicular torque inside the circular head, which results in spiral patterns as shown in (c), also increasing the spread of the fragments. In contrast, the “top-down” view for the rectangular shatter head in (c) indicates, that the large main fragments leave the shatter head centrally on the “shatter plane” (top side) and only the tiny fragments and dust are pushed towards and around the corners.
@@
Following extensive laboratory commissioning and the analysis of fast camera recordings of the resulting pellet fragment clouds [27, 28], three different shatter heads (all with miter bends) were selected for the 2022 experimental campaign [22] (see figure 1(a, b) and table 1). A short, circular cross-section, $25^{\circ}$ head is used for increased spatial spread of the fragments. Two long, rectangular cross-section heads were installed for better collimation. These are a “matching pair” with shatter angles of $12.5^{\circ}$ and $25^{\circ}$ respectively, which allows the more convenient matching of the normal impact velocity ($\text{v}_{\perp}$) – the main factor for the pellet fragment size distribution – at different parallel penetration speeds, as $\sin(25^{\circ})\approx 2\cdot\sin(12.5^{\circ})$. Hereby, a higher normal impact velocity will typically result in smaller fragments [29, 30, 27]. This allows to test the effects of the fragment size distributions and injection velocities on the disruptions independently.
@@
Figure 2 shows typical spatial distributions of the fragments (“side view”) for the $25^{\circ}$, circular, and shortened (a, c) and $12.5^{\circ}$, rectangular, and long (b, e) shatter head as observed during the commissioning phase. Overall the rectangular shatter head has a more collimated fragment plume – with fragments leaving the top side (“shatter plane”) centrally and only the small fragments pushed towards and around the corners (compare figure 2(b)) as shown by the “top-down” video recordings (see figure 2(e)). The circular shatter head was selected for its increased spatial spread of the fragments, however, comes with a larger uncertainty in the fragment distributions. The effective shatter angle $\beta$ is a function of the impact position of the pellet inside the tube (defined via the angle $\theta$ indicated in figure 2(d)). Therefore a larger statistical variance was observed for the circular shatter head in the laboratory commissioning phase [27]. Additionally, the fragments are able to rotate inside the shatter head after impact, which can lead to strong spiral patterns in the resulting fragment plume as shown in figure 2(c). More details on fragment size and velocity distributions are available in the thesis by T. Peherstorfer [27]. A fragment segmentation and analysis pipeline based on machine learning is also currently under development [28]. Different barrels – installed ahead of the experiments – allow to freeze pellets with different diameters. In the 2022 campaign, we used $3\times 4$ mm and $3\times 8$ mm diameter barrel setups. The nominal maximum pellet length over diameter (L/D) of the pellets is 1.7 (4 mm) and 1.2 (8 mm). The pellets can be shortened with the help of the barrel heating (BH) coils [24]. In the 2022 campaign a total of $\sim$240 dedicated SPI discharges were executed.
@@ 2.2 Radiated power measurements
with the power measurement $\text{P}_{\text{ch}}$ for each individual channel depicted in figure 1(a). The weighting coefficients $w_{\text{ch}}=[2.18/16,\ 2.54/16,\ 2.41/16,\ 1.74/16]$ for each of the 4 channels were derived taking into account the geometric effects of the camera. The observation volumes for the four channels shown in figure 1(a) (full-shadow volumes) together do not cover the entire volume of the plasma inside each of the five measurement sectors. Therefore, weighting factors are introduced to calculate the total radiated power for each of the measurement sectors and the entire plasma from the measurement in just a single sector by assuming a uniform radiation (toroidally symmetric) inside the chamber and taking into account geometrical effects – namely the OV-chamber-ratio and half/full-shadow (on the sensor) effects of the pinholes to create the channel. Instead of comparing the calculated total radiated power as derived by the 5 measurements with each other, by dividing these weighting factors by 16 (number of sectors in AUG see figure 1(c)), the radiated power for each individual measurement sector (instead of the entire torus) is derived by the weighted sum of the channels in that particular sector as given in equation (1). For the entire calculation methodology described in this section, we assume an uncertainty of about 20% comparing it to other measurement techniques. However, for a shot-to-shot comparison of the radiated energy values for the SPI experiments, we assume a measurement uncertainty in the range of 10% as they are determined in the same way and no degradation of the sensitivity of the foil bolometers has been observed over the entire shot range.
@@ 3 Definition and calculation of the radiated energy fraction $\text{f}_{\text{rad}}$
The fraction of total plasma energy radiated during the entired disruption – the radiated energy fraction ($\text{f}_{\text{rad}}$) – is calculated following the formula by Lehnen et al [35] and Sheikh et al [36]:
@@
with the additional $\text{W}_{\text{heat}}$ term, representing the plasma heating after the injection. The fixed pre-injection values of $\text{W}_{\text{mag}}$ (magnetic) and $\text{W}_{\text{th}}$ (thermal) are used, while $\text{W}_{\text{rad}}$ (radiated) and $\text{W}_{\text{heat}}$ are calculated from the start of the injection until the end-marker, which will lead to $\text{f}_{\text{rad}}$ (t) as a function of time, which is evaluated at the end of the integration interval (discussed in the following section 3.1). In the following, the methods to calculate the elements of equation (2) are presented.
@@ 3.1 Estimation of the radiated energy $\text{W}_{\text{rad}}$
Prior to the first, large fragments entering the plasma, a slight radiation increase in the outermost line of sight (LOS) (directed towards the LFS midplane) of the fast AXUV camera [31, 32] in sector 16 is observed, which is referred to as the First Light (FL), and is used as the marker for the beginning of the injection phase. This is assumed to be caused by the fragments or small amounts of gas (produced during the pellet break-up) entering the plasma followed by the bulk of the fragments shortly afterwards. The term $\text{W}_{\text{rad}}$ is the estimate for the energy radiated from the entire plasma measured via the foil bolometers, where linear interpolations of sectors with no measurements are used. As a start of the integration interval the FL-marker ($\text{t}_{\text{FL}}$) is used. The end-marker ($\text{t}_{\text{end}}$) is provided by the end of the radiation peak(s), when the derivative of the radiation approaches zero again. The methodology used to derive this end-marker is described in the paper by P. Heinrich et al [37] where it is applied to the plasma current signal to detect the end of the CQ. For disruptive SPI, this end-marker is usually at the end of the CQ ($\text{t}_{\text{CQ-end}}$). For non-disruptive shattered pellet injections, the end-marker is usually around the minimum of the plasma current, before the plasma current starts to recover to its pre-injection level.
The radiated power is calculated for each individual sector, where for sectors without direct measurements, a linear interpolation between the measurement sectors was chosen. Finally, the total radiated power is given by the sum over all 16 sectors:

### v2

@@ Abstract
Future large tokamaks will operate at high plasma currents and high stored plasma energies. To ensure machine protection in case of a sudden loss of plasma confinement (major disruption), a large fraction of the magnetic and thermal energy must be radiated to reduce thermal loads. The disruption mitigation system for ITER is based on massive material injection in the form of shattered pellet injection (SPI). To support ITER, a versatile SPI system was installed at the tokamak ASDEX Upgrade (AUG). The AUG SPI features three independent pellet generation cells and guide tubes, and each was equipped with different shatter heads for the 2022 experimental campaign. We dedicated over 200 plasma discharges to the study of SPI plasma termination, and in this manuscript report on the results of bolometry (total radiation) analysis. We found, that the amount of neon inside the pellets is the dominant factor determining the radiated energy fraction ($\text{f}_{\text{rad}}$). Large and fast fragments, produced by the $12.5^{\circ}$ rectangular shatter head, lead to somewhat higher values of $\text{f}_{\text{rad}}$ compared to the $25^{\circ}$ circular or rectangular heads. This effect is strongest for neon content of $\lesssim 3\times 10^{20}$ neon atoms ($\textrm{f}_{\textrm{neon}}\lesssim 1.25\%$ neon) injected, where a lower normal velocity component (larger fragments) seems slightly beneficial. While full-sized, 8 mm diameter, 100% deuterium ($\textrm{D}_{2}$) pellets lead to a disruption, the 4 mm or shortened 8 mm pellets of 100% $\textrm{D}_{2}$ did not. The disruption threshold for 100% $\textrm{D}_{2}$ is found to be around $1\times 10^{22}$ deuterium molecules inside the pellet. While the radiated energy fraction of non-disruptive SPI is below 20%, this is increased to 40% during the TQ and VDE phase of the disruptive injections. For deuterium–neon–mix pellets, $\text{f}_{\text{rad}}$-values of $\leq 90$% are observed, and the curve saturates around 80% already for 10% neon mixed into the 8 mm pellets ($2\times 10^{21}$ neon atoms).
@@ 1 Introduction
Plasma terminating disruptions pose a major challenge for tokamaks with high plasma current and stored energy [1, 2, 3]. An efficient disruption mitigation system (DMS) should address three main challenges: minimise the thermal loads on the plasma facing components (PFCs), minimise eddy and halo currents which cause large mechanical loads, while suppressing the formation of a large runaway electron (RE) beam [1]. In present day tokamaks, disruption mitigation is typically realised using massive material injection (MMI), either via massive gas injection (MGI) or shattered pellet injection (SPI). High-Z material inside the injection mix helps spread the energy over a larger surface area via radiation, reducing the localised thermal loads [4, 5]. The mechanical loads can be reduced by tailoring the duration of the current quench (CQ) through the injected material composition [4, 6] or early injections in cases of vertical displacement events (VDEs) [7, 8]. The formation of a RE beam is hoped to be suppressed by increasing the free electron density in the plasma core [9], hence the material assimilation (of low Z material – namely protium/deuterium) is a key factor [10, 11, 12].
The SPI technique was first tested at DIII-D in 2009 [13, 14], and over time more tokamaks, such as JET [14], KSTAR [15, 16], HL-2A [17], J-TEXT [18, 19], and EAST [20, 21] followed.
In late 2021, a highly flexible SPI system was installed at ASDEX Upgrade (AUG) [22, 23, 24, 25] to provide further input for the design and optimisation of the ITER DMS [26, 10] based on the SPI principle [27]. In this paper we discuss the evolution of total radiated energy ($\text{W}_{\text{rad}}$) and the radiated energy fraction ($\text{f}_{\text{rad}}$) as a function of pellet and injection parameters in different plasma scenarios, based on the bolometry analysis of the 2022 AUG SPI experimental campaign.
@@ 2.1 Shattered pellet injector on AUG
The AUG SPI is a triple-barrel system, with three independent guide tubes and three different shatter heads [22, 24]. Three pellets, made of mixtures of deuterium and neon (as well as 100% deuterium or neon) can be generated and launched simultaneously, potentially allowing the study of multi-injection scenarios. A detailed description of the SPI system is given in the paper by Dibon et al [22] and the laboratory commissioning [^2] Animation video of the laboratory setup available at https://datashare.mpcdf.mpg.de/s/DlMzGcWnZwoHMjq or [24]. by Heinrich et al [24, 25].
@@
Figure 2: Comparison of the spatial distribution of the fragment plumes generated by different shatter heads. In (a) a larger spread of the fragments is observed for the circular, $25^{\circ}$ shatter head compared to the more collimated fragment plume of the rectangular, $12.5^{\circ}$ head for similar pre-shattering pellet velocity in (b). In (c) and (d) the geometrical effects of the circular shatter head are shown. The effective shatter angle $\beta$ depends on the impact position ($\theta$) inside a circular shatter head [28], while for the rectangular head the shatter angle is equal to the miter bend angle $\alpha$. Additionally, the shards may receive perpendicular torque inside the circular head, which results in spiral patterns as shown in (c, d), also increasing the spread of the fragments. After leaving the exit of the shatter head, the spatial distribution is set by the position $\varphi$ at the time of exit, depending on the rotation inside the head. This is indicated by the red, dashed line in (c, d). In contrast, the “top-down” view for the rectangular shatter head in (e) indicates, that the large main fragments leave the shatter head centrally on the “shatter plane” ($\theta=\varphi\approx 0^{\circ}$; top side) and only the tiny fragments and dust are pushed towards and around the corners ($\varphi>0^{\circ}$). Hereby the fragments seem to have a minor triangular fragment plume, similar to the observations by T. Gebhart et al [32]. In the side view in (b), this will result in the collimated, arched fragment plume: large fragments will have a minor displacement to the shatter plane, whereas small fragments will be pushed around the corners and appear further downwards (with larger displacement).
@@
Following extensive laboratory commissioning and the analysis of fast camera recordings of the resulting pellet fragment clouds [28, 29], three different shatter heads (all with miter bends) were selected for the 2022 experimental campaign [22] (see figure 1(a, b) and table 1). A short, circular cross-section, $25^{\circ}$ head is used for increased spatial spread of the fragments. Two long, rectangular cross-section heads were installed for better collimation. These are a “matching pair” with shatter angles of $12.5^{\circ}$ and $25^{\circ}$ respectively, which allows the more convenient matching of the normal impact velocity ($\text{v}_{\perp}$) – the main factor for the pellet fragment size distribution. As $\sin(25^{\circ})\approx 2\cdot\sin(12.5^{\circ})$, different parallel velocity components – hence fragment velocities [28] – are achieved while matching $\text{v}_{\perp}$. Hereby, a higher normal impact velocity will typically result in smaller fragments [30, 31, 28, 32]. This allows to test the effects of the fragment size distributions and injection velocities on the disruptions independently.
@@
Figure 2 shows typical spatial distributions of the fragments (“side view”) for the $25^{\circ}$, circular, and shortened (a, c) and $12.5^{\circ}$, rectangular, and long (b, e) shatter head as observed during the commissioning phase. Overall the rectangular shatter head has a more collimated fragment plume – with fragments leaving the top side (“shatter plane”) centrally and only the small fragments pushed towards and around the corners (compare figure 2(b)) as shown by the “top-down” video recordings (see figure 2(e)). This will create an arched fragment plume as the large fragments leaving the head at the shatter plane show a low displacement w.r.t. the upper edge of the shatter head compared to the smaller fragments, that are pushed around the corners, hence showing a larger displacement. For the top-down view, the fragment plume of the rectangular head seems more spread out, with a slight triangular pattern indicated in figure 2(e) as also observed by T. Gebhart et al [32]. Compared to the so-called “horn” geometry (rectangular head with fixed height but widening width [25]), the only difference is observed for the tiny fragments and dust, pushed all the way to the edges, thereby producing less of a pronounced “dust cloud”. The large fragments behave exactly like in the classical rectangular geometry in figure 2(e) and do not spread alongside the widening shatter plane. The circular shatter head was selected for its increased spatial spread of the fragments (in all directions), however, comes with a larger uncertainty in the fragment distributions. The effective shatter angle $\beta$ is a function of the impact position of the pellet inside the tube (defined via the angle $\theta$ indicated in figure 2(d)). Therefore a larger statistical variance was observed for the circular shatter head in the laboratory commissioning phase [28]. Additionally, the fragments are able to rotate inside the shatter head after impact, which can lead to strong spiral patterns in the resulting fragment plume as shown in figure 2(c). While the motion of the fragments is linear after exiting the shatter head, the spatial distribution – defined by the angle $\varphi$ at the exit – depends on this spiral motion and larger fragments are not only observed close to the upper edge of the exit as illustrated for the rectangular geometry in figure 2(b, e). More details on fragment size and velocity distributions are available in the thesis by T. Peherstorfer [28]. A fragment segmentation and analysis pipeline based on machine learning is also currently under development [29]. Different barrels – installed ahead of the experiments – allow to freeze pellets with different diameters. For the 2022 campaign, we installed a set of three 4 mm diameter barrels at first, and during the campaign switched to a set of three 8 mm diameter barrels. The nominal maximum pellet length over diameter (L/D) of the pellets is 1.7 (4 mm) and 1.2 (8 mm) [24]. The pellets can be shortened with the help of the barrel heating (BH) coils. In the 2022 campaign a total of $\sim$240 dedicated SPI discharges were executed.
@@ 2.2 Radiated power measurements
$\text{W}_{\text{plasma}}$ calculated eq. (2) energy dissipated in the plasma [37]
@@
with the power measurement $\text{P}_{\text{ch}}$ for each individual channel depicted in figure 1(a). The weighting coefficients $w_{\text{ch}}=[2.18/16,\ 2.54/16,\ 2.41/16,\ 1.74/16]$ for each of the 4 channels were derived taking into account the geometric effects of the camera. The observation volumes for the four channels shown in figure 1(a) (full-shadow volumes) together do not cover the entire volume of the plasma inside each of the five measurement sectors. Therefore, weighting factors are introduced to calculate the total radiated power for each of the measurement sectors or even the entire plasma from the measurement in just a single sector. Hereby, a toroidally uniform radiation is assumed within the sector ($\varphi=22.5^{\circ}$) or the entire plasma ($\varphi=360^{\circ}$) and geometrical effects are taken into account – namely the half/full-shadow (on the sensor) effects of the pinholes to create the channel. Within this paper, we use the weighting factors divided by 16 (number of sectors in AUG see figure 1(c)) to calculate the radiated power for each individual measurement sector (instead of the entire torus) as given in equation (1). More details on the derivations are available in the PhD thesis by P. Heinrich [25].
For the entire calculation methodology described in this section, we assume an uncertainty of about 20% comparing it to other measurement techniques. However, for a shot-to-shot comparison of the radiated energy values for the SPI experiments, we assume a measurement uncertainty in the range of 10% as they are determined in the same way and no degradation of the sensitivity of the foil bolometers has been observed over the entire shot range.
@@ 3 Definition and calculation of the radiated energy fraction $\text{f}_{\text{rad}}$
The fraction of total plasma energy radiated during the entire disruption – the radiated energy fraction ($\text{f}_{\text{rad}}$) – is calculated following the formula by Lehnen et al [37] and Sheikh et al [38]:
@@
with the additional $\text{W}_{\text{heat}}$ term, representing the plasma heating after the injection. The fixed pre-injection values of $\text{W}_{\text{mag}}$ (magnetic) and $\text{W}_{\text{th}}$ (thermal) are used, while $\text{W}_{\text{rad}}$ (radiated) and $\text{W}_{\text{heat}}$ are calculated from the start of the injection until the end-marker, which will lead to $\textrm{f}_{\textrm{rad}}$(t) as a function of time, which is evaluated at the end of the integration interval (discussed in the following section 3.1).
Figure 3: Energy balance during the disruption. The magnetic and thermal energy of the plasma are converted into $\text{W}_{\text{c}}\ (+\textrm{W}_{\textrm{RE}})+\text{W}_{\text{rad}}+\textrm{W}_{\textrm{con}}$, with $\textrm{W}_{\textrm{con}}$ being the conducted and convected energy to the plasma facing components. In the AUG SPI experiments, the population of runaway electrons (REs) is negligible, hence the term $\textrm{W}_{\textrm{RE}}$ is set to zero for the analysis. The term $\text{W}_{\text{heat}}$ describes the external plasma heating after the injection (see section 3.3). The energy dissipated in the plasma ($\text{W}_{\text{plasma}}$) after the disruption is therefore given as $\text{W}_{\text{mag}}+\text{W}_{\text{th}}+\text{W}_{\text{heat}}-\text{W}_{\text{c}}$.
In the following, the methods to calculate the elements of equation (2) are presented.
@@ 3.1 Estimation of the radiated energy $\text{W}_{\text{rad}}$
Prior to the first, large fragments entering the plasma, a slight radiation increase in the outermost line of sight (LOS) (directed towards the low-field side (LFS) midplane) of the fast AXUV camera [33, 34] in sector 16 is observed, which is referred to as the First Light (FL), and is used as the marker for the beginning of the injection phase. This is assumed to be caused by the fragments or small amounts of gas (produced during the pellet break-up) entering the plasma followed by the bulk of the fragments shortly afterwards. The term $\text{W}_{\text{rad}}$ is the estimate for the energy radiated from the entire plasma measured via the foil bolometers, where linear interpolations of sectors with no measurements are used. As a start of the integration interval the FL-marker ($\text{t}_{\text{FL}}$) is used. The end-marker ($\text{t}_{\text{end}}$) is provided by the end of the radiation peak(s), when the derivative of the radiation approaches zero again. The methodology used to derive this end-marker is described in the paper by P. Heinrich et al [39] where it is applied to the plasma current signal to detect the end of the CQ. For disruptive SPI, this end-marker is usually at the end of the CQ ($\text{t}_{\text{CQ-end}}$). For a non-disruptive injection, the end-marker is usually around the minimum of the plasma current, before the plasma current starts to recover to its pre-injection level.
The radiated power is calculated for each individual sector, where for sectors without direct measurements, a linear interpolation between the measurement sectors was chosen. Here, the highest peaking is observed for low neon or 100% $\textrm{D}_{2}$ injections in sector 16 (injection location), where the neighboring sectors 1 and 15 already show radiation levels very close to the sectors 5 and 9, hence, a linear interpolation was suggested [40]. Finally, the total radiated power is given by the sum over all 16 sectors: